# Choose Your Battles: Distributed Learning Over Multiple Tug of War Games

### v1

@@ Abstract
Consider $N$ players and $K$ games taking place simultaneously. Each of these games is modeled as a Tug-of-War (ToW) game where increasing the action of one player decreases the reward for all other players. Each player participates in only one game at any given time. At each time step, a player decides the game in which they wish to participate in and the action they take in that game. Their reward depends on the actions of all players that are in the same game. This system of $K$ games is termed ‘Meta Tug-of-War’ (Meta-ToW) game. These games can model scenarios such as power control, distributed task allocation, and activation in sensor networks. We propose the Meta Tug-of-Peace algorithm, a distributed algorithm where the action updates are done using a simple stochastic approximation algorithm, and the decision to switch games is made using an infrequent 1-bit communication between the players. We prove that in Meta-ToW games, our algorithm converges to an equilibrium that satisfies a target Quality of Service reward vector for the players. We then demonstrate the efficacy of our algorithm through simulations for the scenarios mentioned above.
@@ 1 INTRODUCTION
In many network scenarios, a conflict arises between the agents regarding how to share resources and tasks. Examples are transmitter power control in wireless networks, activation probability in sensor networks, and multi-robot workers. Each agent has a local performance measure, such as its throughput, energy consumption, or progress rate. When a wireless device uses more power to transmit, it increases its throughput but decreases that of others. When a sensor, which collects its own data but also relays that of others, is activated with lower probability, it improves its energy consumption but worsens the packet loss rate of others. When a robot spends effort on a task, it increases its progress rate but decreases the marginal contribution of others, possibly pushing them to switch to a task where they can be more efficient. These conflicts can naturally be modeled as a game between the agents.
A conflict between agents, however, does not mean that agents are selfish. The agents are programmable and could run the protocol we design (e.g., wireless protocols, such as WiFi or 5G). Moreover, the agents (or players) often have a minimum requirement of reward, which we term as their Quality of Service (QoS). We wish to design a distributed algorithm that allows each player to achieve their QoS asymptotically. The challenge in designing such a protocol is that players often just observe stochastic bandit feedback, i.e., a noisy version of the reward at the current action profile. The players do not know their reward function, the actions of other players, or the reward received by other players. Hence the challenge is to know whether to “insist” on improving their local reward or give up in favor of others. In large-scale networks with thousands or millions of devices, this coordination challenge is the primary design bottleneck.
@@
We first design the Tug-of-Peace algorithm (Algorithm 1), a simple distributed algorithm for a ToW game. This algorithm converges with probability 1 to an action profile that satisfies the given QoS requirements for all players, if the QoS requirements are feasible. Additionally, we show that with high probability, Algorithm 1 converges to the minimal action profile (Definition 3) which satisfies the QoS requirements. This algorithm requires infrequent (i.e., at finite timesteps), one-bit communications. We also present an algorithm (Algorithm 2) for the case where even one-bit communication is not possible. We show that this algorithm also converges to the minimal action profile with high probability. We name these algorithms Tug-of-Peace (ToP) since, instead of pulling the rope as hard as they can, players learn to place the marker at a point where they can all achieve their target QoS.
In the meta game, the players need to find a configuration of players to different games such that the QoS requirements are feasible in that configuration. We design Meta-ToP algorithm (Algorithm 3), a distributed algorithm in which players use the infrequent one-bit communication to communicate to other players that the QoS requirement ‘may’ not be feasible in the current configuration of players. On receiving these communications, the players switch to different games, allowing players to try different configurations. We show that our algorithm attains a configuration and converges to an action profile that satisfies the QoS requirements for all players with probability 1. Moreover, we show that the game switches happen only a finite number of times.
Finally, we demonstrate in simulations the performance of our algorithms for power control, distributed task allocation and sensor activation.
@@ 1.1 Related Work
QoS is a common objective in transmission power control in wireless networks [1, 2, 3, 4, 5]. This application is a special case of our model, which we study in Subsection 5.1. The algorithms presented by these works for QoS in power control exploit the form of the reward function, which they assume is known. Our algorithm, on the other hand, does not need to know the reward functions, as long as the game is a ToW game.
@@
The analysis methods of [1, 2, 3, 4] are specific to the power control wireless scenario. Their analysis depends on the eigenvalues of a matrix defined by using the channel gain matrix and the QoS vector. As an example, [2] models the problem as finding the solutions of a linear system, and its analysis then depends on the eigenvalue properties of this linear system. This model and properties are specific to that scenario and cannot be extended to our broader class of games. Other works such as [5] take a general approach, but also with certain limitations. They use the properties of monotone and sub-homogeneous maps to show the convergence of the algorithm they design. These properties are facilitated by the structure of power control games. However, designing an algorithm that satisfies these properties is challenging for other general games. Additionally, our algorithm is better equipped to deal with noise than the algorithm by [5], as we show later.
@@
Beyond power control, distributed protocols with QoS guarantees were studied in [6, 7]. The work in [6] considered a multiplayer bandit scenario where, if multiple players pick the same arm, they all receive zero reward. This is a special case of a game that is different from the class of games we consider. The work in [7] considers a general discrete game, but the algorithm uses regular communication between the players. Compared to both algorithms in [6, 7], our proposed algorithm here has substantially better scalability in the number of players. Furthermore, our algorithm has two variants that either require 1-bit of infrequent communication or no communication at all. QoS guarantees were also studied in [8], but in the context of average cost setting in a multi-agent Markov decision process.
Our work is also related to the growing literature on open multi-agent systems, and open games in particular, where players can join and leave the game [9, 10].
To analyze the convergence of our algorithm, we use the Ordinary Differential Equation (ODE) approach to stochastic approximation [11] and the concept of Cooperative ODEs and Monotone Dynamical Systems [12]. Apart from their use in dynamical systems, cooperative ODEs have been widely used in fields such as epidemiology [13, 14], social networks [15, 16], and queuing systems [17].
@@ 1.2 Notation
We use bold letters to denote vectors and $\mathbf{0}_{N}$ and $\boldsymbol{\delta}_{N}$ to denote the N-dimensional vector with all zeros and all $\delta$, respectively. We also use $\Pi_{\mathcal{X}}$ to denote the Euclidean projection into the set $\mathcal{X}$ and $\|\cdot\|$ to denote the Euclidean norm. We use the standard vector inequalities, where $\boldsymbol{x}\leq\boldsymbol{y}$ denotes $x_{i}\leq y_{i}$ for all $i$. Similarly, we use the notation $\boldsymbol{x}\in[\boldsymbol{z},\boldsymbol{y}]$ to denote $z_{i}\leq x_{i}\leq y_{i}$ for all $i$.
We use bold letters to denote vectors and matrices, and the standard game-theoretic notation where $\boldsymbol{x}_{-n}$ is the vector of actions for all players except player $n$. $\|\cdot\|$ denotes the $\ell_{2}$-norm throughout the paper when applied on a vector, and the operator norm for matrices (under the $\ell_{2}$-norm for vectors).
@@ 2 PROBLEM FORMULATION
Consider a set of $N$ players $\mathcal{N}=\{1,\ldots,N\}$ and a set of $K$ games $\mathcal{G}=\{1,\ldots,K\}$. Each player $n$ decides the game $g_{n}\in\mathcal{G}$ they wish to participate in and the action $x_{n}\in\mathcal{X}_{n}$ they take in that game. Here $\mathcal{X}_{n}\coloneqq[0,B_{n}]\subseteq{\mathbb{R}^{+}}$. Let $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{N}$. Also define the interior of this region as $\mathcal{X}^{\textrm{o}}=(0,B_{1})\times\ldots\times(0,B_{N})$. Each player $n$ has a reward function $u_{n}(\boldsymbol{g},\boldsymbol{x})$ that depends only on the actions of the players in the game $g_{n}$. Here $\boldsymbol{g}\in\mathcal{G}^{N}$ denotes the game configuration of players, or the vector of the games each player is participating in, i.e., $\boldsymbol{g}=[g_{1},\ldots,g_{N}]$ and $\boldsymbol{x}$ is the action profile for all players, i.e., $\boldsymbol{x}=[x_{1},\ldots,x_{N}]$.
Before defining general Meta-ToW games, we define a single ToW game. As there is only a single game, we use $u_{n}(\boldsymbol{x})$ to denote the reward for player $n$.
@@
The reward function also satisfies $u_{n}(\boldsymbol{g},\boldsymbol{x})=0$, if $x_{n}=0$ and $u_{n}(\boldsymbol{g},\boldsymbol{x})\geq 0$ for all $n,\boldsymbol{x}\in\mathcal{X}$.
Each player chooses the game $g_{n}(t)$ and their action $x_{n}(t)$ at each timestep $t\geq 0$, and receives noisy reward $y_{n}(t)=u_{n}(\boldsymbol{g}(t),\boldsymbol{x}(t))+M_{n}(t)$. The players receive no information about other players’ games, actions, and rewards. Here $\mathbf{M}(t)=[M_{1}(t),\ldots,M_{N}(t)]$ denotes the noise sequence and satisfies the following assumption.
Assumption 1.
$\mathbf{M}(t)$ is a Martingale difference sequence, i.e., $\mathbb{E}[\mathbf{M}(t)|\mathcal{F}_{t-1}]=0$ where $\mathcal{F}_{t}\coloneqq\sigma(\boldsymbol{x}(s),\mathbf{M}(s),s\leq t)$. Additionally, $\mathbf{M}(t)$ has bounded support, i.e., $|M_{n}(t)|\leq\widehat{M}$, w.p. $1$, for all $n\in\mathcal{N},t\geq 0$ for some positive $\widehat{M}$.
Our aim is to design a distributed algorithm such that each player $n$ asymptotically achieves a reward of at least $\lambda_{n}>0$, which we collectively denote as the Quality of Service (QoS) vector $\boldsymbol{\lambda}$. This is not possible for all $\boldsymbol{\lambda}$ and hence we make the following feasibility assumption.
Assumption 2.
We make the following assumptions.
  1. (a)
    There exists $\boldsymbol{x}\in\mathcal{X}^{\textrm{o}},\boldsymbol{g}\in\mathcal{G}^{N}$ and $\delta>0$ such that $u_{n}(\boldsymbol{g},\boldsymbol{x})\geq\lambda_{n}+\delta$ for all players $n$.
  2. (b)
    Let $\boldsymbol{x}_{g}$ denote the vector of players’ numerical action whose game is $g$ and $Du(\boldsymbol{x}_{g})$ denote the Jacobian at $\boldsymbol{x}_{g}$, i.e., the Jacobian for players’ actions whose game is $g$. Suppose $\boldsymbol{\overline{\lambda}}$ is chosen uniformly at random in $\boldsymbol{\lambda},\boldsymbol{\lambda}+\boldsymbol{\delta}_{N}]$. If $u(\hat{\boldsymbol{g}},\hat{\boldsymbol{x}})=\boldsymbol{\overline{\lambda}}$, then w.p. 1 the Jacobian $Du(\hat{\boldsymbol{x}}_{g})$ has no purely imaginary eigenvalues for all $g$, where the game configuration is given by vector $\hat{\boldsymbol{g}}$.
The first part of this assumption ensures that the QoS vector of rewards is feasible and can be achieved in the interior of the set $\mathcal{X}$. Here $\delta$ can be infinitesimally small. There could be multiple such points. Since the action of the player is a measure of “pull” (e.g., power), we prefer equilibrium points with smaller actions.
The parameter $\delta$ is also needed for technical reasons. For an arbitrary QoS vector $\boldsymbol{\lambda}$, the QoS might be achieved at an equilibrium point that is not stable for our algorithm [21]. Under Assumption 2, the equilibrium point is stable with probability 1 if we take a QoS vector at random in $\boldsymbol{\lambda},\boldsymbol{\lambda}+\boldsymbol{\delta}_{N}]$.
The second part of the assumption is very mild for matrices with non-zero eigenvalues, since perturbing a matrix that has purely imaginary eigenvalues adds a non zero real part to these eigenvalues[17], [21, Section 8.4]. Hence, for randomly chosen $\boldsymbol{\overline{\lambda}}$, the Jacobian $Du(\hat{\boldsymbol{x}}_{g})$ does not have purely imaginary eigenvalues with probability 1.

### v2

@@ Abstract
Consider $N$ players and $K$ games taking place simultaneously. Each of these games is modeled as a Tug-of-War (ToW) game where increasing the action of one player decreases the reward for all other players. Each player participates in only one game at any given time. At each time step, a player decides the game in which they wish to participate in and the action they take in that game. Their reward depends on the actions of all players that are in the same game. This system of $K$ games is termed a ‘Meta Tug-of-War’ (Meta-ToW) game. These games can model scenarios such as power control, distributed task allocation, and activation in sensor networks. We propose the Meta Tug-of-Peace algorithm, a distributed algorithm where the action updates are done using a simple stochastic approximation algorithm, and the decision to switch games is made using an infrequent 1-bit communication between the players. We prove that in Meta-ToW games, our algorithm converges to an equilibrium that satisfies a target Quality of Service reward vector for the players. We then demonstrate the efficacy of our algorithm through simulations for the scenarios mentioned above.
@@ 1 INTRODUCTION
In many network scenarios, a conflict arises between the agents regarding how to share resources and tasks. Examples are transmitter power control in wireless networks, activation probability in sensor networks, and task allocation in multi-robot networks. Each agent has a local performance measure, such as its throughput, energy consumption, or progress rate. When a wireless device uses more power to transmit, it increases its throughput but decreases that of others. When a sensor, which collects its own data but also relays that of others, is activated with a lower probability, it improves its energy consumption but worsens the packet loss rate of others. When a robot spends effort on a task, it increases its progress rate but decreases the marginal contribution of others, possibly pushing them to switch to a task where they can be more efficient. These conflicts can naturally be modeled as a game between the agents.
A conflict between agents, however, does not mean that agents are selfish. The agents are programmable and run the protocol we design (e.g., WiFi or 5G). Moreover, the agents (or players) often have a minimum requirement of reward, which we term as their Quality of Service (QoS).
Finding an action profile that satisfies the QoS guarantees is an optimization problem that can be solved by a centralized server that instructs the agents how to play. However, such a centralized server would have to know the reward functions and action sets, which implies that agents must send this information to the server. The server then needs to compute the solution in real-time and send it back to the agents. This creates latencies and limits the system’s scalability, since the complexity of the optimization problem grows fast with the number of agents. Such a centralized scheme also violates user privacy and creates a security vulnerability, since if the server is hacked, the hacker gains control of all agents. Therefore, distributed protocols are preferred [15, 40, 4, 36, 38, 29, 31].
We wish to design a distributed algorithm that allows each player to achieve their QoS asymptotically. The challenge in designing such a protocol is that players often just observe stochastic bandit feedback, i.e., a noisy version of the reward at the current action profile. The players do not know their reward functions and do not observe the actions or rewards of others. Hence, the challenge is to know whether to “insist” on improving their local reward or give up in favor of others. In large-scale networks with thousands or millions of devices, this coordination challenge is the primary design bottleneck.
@@
We first design the Tug-of-Peace algorithm (Algorithm 1), a simple distributed algorithm for a ToW game. This algorithm converges with probability 1 to an action profile that satisfies the given QoS requirements for all players, if the QoS requirements are feasible. Additionally, we show that with high probability, Algorithm 1 converges to the minimal action profile (Definition 3) which satisfies the QoS requirements. This algorithm requires infrequent (i.e., finitely many), one-bit communications. We also present an algorithm (Algorithm 2) for the case where even one-bit communication is not possible. We show that this algorithm also converges to the minimal action profile with high probability. We name these algorithms Tug-of-Peace (ToP) since, instead of pulling the rope as hard as they can, players learn to place the marker at a point where they can all achieve their target QoS.
In the meta game, the players need to find a configuration of players to different games such that the QoS requirements are feasible in that configuration. We design the Meta-ToP algorithm (Algorithm 3), a distributed algorithm in which players use the infrequent one-bit communication to communicate to other players that the QoS requirement ‘may’ not be feasible in the current configuration of players. On receiving these communications, the players switch to different games, allowing players to try different configurations. We show that our algorithm attains a configuration and converges to an action profile that satisfies the QoS requirements for all players with probability 1. Moreover, we show that the game switches happen only a finite number of times. Finally, we demonstrate in simulations the performance of our algorithms for power control, distributed task allocation and sensor activation.
@@ 1.1 Related Work
Distributed algorithms that converge to a Nash equilibrium (NE) have been extensively studied for various games [39, 17, 27, 16, 34]. The NE can be globally inefficient, and in particular, does not satisfy any QoS guarantees. Instead of a NE, the equilibrium our players seek in this paper is an action profile where all of their rewards are above the QoS threshold.
Distributed protocols with QoS guarantees for general games are much less studied compared to distributed algorithms that converge to NE. The work in [5] considered a multiplayer bandit scenario where, if multiple players pick the same arm, they all receive zero reward. This is a special case of a game that is different from the class of games we consider. The work in [6] considers a general discrete game, but the algorithm uses regular communication between the players. Compared to both algorithms in [5, 6], our proposed algorithm here has substantially better scalability in the number of players. Furthermore, our algorithm has two variants that either require 1-bit of infrequent communication or no communication at all. QoS guarantees were also studied in [24], but in the context of average cost setting in a multi-agent Markov decision process.
Distributed protocols with QoS guarantees are more common for power control in wireless networks, specifically [15, 40, 4, 36, 38]. The power control game is a special case of the Meta-ToW games we consider in this paper, as explained in Subsection 3.1. The power control QoS algorithms exploit the form of the reward function, which they assume is known. Our algorithm, on the other hand, does not need to know the reward functions as long as the game is a ToW game.
@@
The analysis methods of [15, 40, 4, 36] are specific to the power control wireless scenario. Their analysis depends on the eigenvalues of a matrix defined using the channel gain matrix and the QoS vector. As an example, [40] models the problem as finding the solutions of a linear system, and its analysis then depends on the eigenvalue properties of this linear system. This model and properties are specific to that scenario and cannot be extended to our broader class of games. Other works, such as [38], take a general approach, but also with certain limitations. They use the properties of monotone and sub-homogeneous maps to show the convergence of the algorithm they design. These properties are facilitated by the structure of power control games. However, designing an algorithm that satisfies these properties is challenging for other general games. Additionally, our algorithm is better equipped to deal with noise than the algorithm by [38], as we show later.
@@
A Meta-ToW game consists of multiple games that players can join and leave. Thus, our work is related to the growing literature on open multi-agent networks[14, 7, 28].
Our analysis employs the Ordinary Differential Equation (ODE) approach to stochastic approximation [9] and the concept of Cooperative ODEs and Monotone Dynamical Systems [21]. Other than in dynamical systems, cooperative ODEs have been widely used in fields such as epidemiology [26, 18], social networks [3, 2], and queuing systems [11].
@@ 1.2 Notation
We use bold letters to denote vectors and matrices and $\mathbf{0}_{N}$ and $\boldsymbol{\delta}_{N}$ to denote the $N$-dimensional all-zero and all-$\delta$ vectors, respectively. We use $\Pi_{\mathcal{X}}$ to denote the Euclidean projection into the set $\mathcal{X}$. We use the standard vector inequalities, where $\boldsymbol{x}\leq\boldsymbol{y}$ denotes $x_{i}\leq y_{i}$ for all $i$ and $\boldsymbol{x}\in[\boldsymbol{z},\boldsymbol{y}]$ denotes $z_{i}\leq x_{i}\leq y_{i}$ for all $i$. We use the standard game-theoretic notation where $\boldsymbol{x}_{-n}$ is the vector of actions for all players except player $n$. $\|\cdot\|$ denotes the Euclidean norm when applied on a vector, and the operator norm for matrices (under the Euclidean norm for vectors).
@@ 2 Problem Formulation
Consider a set of $N$ players $\mathcal{N}=\{1,\ldots,N\}$ and a set of $K$ games $\mathcal{G}=\{1,\ldots,K\}$. Each player $n$ decides the game $g_{n}\in\mathcal{G}$ they wish to participate in and the action $x_{n}\in\mathcal{X}_{n}$ they take in that game. Here $\mathcal{X}_{n}\coloneqq[0,B_{n}]\subseteq{\mathbb{R}^{+}}$. Let $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{N}$. Define the interior of this region as $\mathcal{X}^{\textrm{o}}=(0,B_{1})\times\ldots\times(0,B_{N})$. Each player $n$ has a reward function $u_{n}(\boldsymbol{g},\boldsymbol{x})$ that depends only on the actions of the players in the game $g_{n}$. Here $\boldsymbol{g}\in\mathcal{G}^{N}$ is the game configuration vector, which indicates the game each player is in, i.e., $\boldsymbol{g}=[g_{1},\ldots,g_{N}]$ and $\boldsymbol{x}$ is the action profile for all players, i.e., $\boldsymbol{x}=[x_{1},\ldots,x_{N}]$.
Define the Quality of Service (QoS) vector $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{N}]$. Our goal is to design a distributed algorithm such that the players, under noisy bandit feedback, converge to a game configuration and action profile such that
$u_{n}(\boldsymbol{g},\boldsymbol{x})\geq\lambda_{n},\quad\forall n\in\mathcal{N}.$ (1)
We next explain the noisy bandit feedback and the required assumptions on the QoS vector.
Each player chooses one game $g_{n}(t)$ to be in, and their action $x_{n}(t)$ in this game, at each timestep $t\geq 0$. The player then receives a noisy reward $y_{n}(t)=u_{n}(\boldsymbol{g}(t),\boldsymbol{x}(t))+M_{n}(t)$. For example, in communication systems, the reward could be the probability to successfully transmit a packet, while the noisy feedback is ACK/NACK. The players receive no information about other players’ games, actions, and rewards. They also cannot observe the rewards they would have received in other games. Here $\mathbf{M}(t)=[M_{1}(t),\ldots,M_{N}(t)]$ denotes the noise sequence and satisfies the following assumption.
Assumption 1.
$\mathbf{M}(t)$ is a martingale difference sequence, i.e., $\mathbb{E}[\mathbf{M}(t)|\mathcal{F}_{t-1}]=0$ where $\mathcal{F}_{t}\coloneqq\sigma(\boldsymbol{x}(s),\boldsymbol{g}(s),\mathbf{M}(s),s\leq t)$. Additionally, $\mathbf{M}(t)$ has bounded support, i.e., $|M_{n}(t)|\leq\widehat{M}$, w.p. $1$, for all $n\in\mathcal{N},t\geq 0$ for some positive $\widehat{M}$.
Not all QoS vectors $\boldsymbol{\lambda}$ are feasible, and hence we make the following feasibility assumption.
Assumption 2.
We make the following assumptions.
  1. (a)
    There exist $\boldsymbol{x}\in\mathcal{X}^{\textrm{o}},\boldsymbol{g}\in\mathcal{G}^{N}$ and $\delta>0$ such that $u_{n}(\boldsymbol{g},\boldsymbol{x})\geq\lambda_{n}+\delta$ for all players $n$.
  2. (b)
    Let $\boldsymbol{x}_{g}$ denote the vector of actions of players that are in game $g$ and $Du(\boldsymbol{x}_{g})$ denote the Jacobian at $\boldsymbol{x}_{g}$, i.e., the Jacobian for players’ actions that are in game $g$. Suppose $\boldsymbol{\overline{\lambda}}$ is chosen uniformly at random in $\boldsymbol{\lambda},\boldsymbol{\lambda}+\boldsymbol{\delta}_{N}]$. If $u(\hat{\boldsymbol{g}},\hat{\boldsymbol{x}})=\boldsymbol{\overline{\lambda}}$, then w.p. 1 the Jacobian $Du(\hat{\boldsymbol{x}}_{g})$ has no purely imaginary eigenvalues for all $g\in\mathcal{G}$, where the game configuration is given by vector $\hat{\boldsymbol{g}}$.
Part (a) of this assumption ensures that the QoS vector is feasible and can be achieved in the interior of the set $\mathcal{X}$. Here $\delta$ can be infinitesimally small.
The parameter $\delta$ is also needed for technical reasons. For an arbitrary QoS vector $\boldsymbol{\lambda}$, the QoS might be achieved at an equilibrium point that is not stable for our algorithm [20]. We later show in Lemma 2 that under Assumption 2, the corresponding equilibrium point is stable with probability $1$ when the QoS vector is drawn uniformly at random from $\boldsymbol{\lambda},\boldsymbol{\lambda}+\boldsymbol{\delta}_{N}]$.
A key requirement for this result is part (b), which is a mild nondegeneracy assumption. Intuitively, this condition is not restrictive: for non-zero purely imaginary eigenvalues, an arbitrarily small perturbation typically introduces a non-zero real part [11], [20, Section 8.4]. Thus, for a randomly chosen $\boldsymbol{\overline{\lambda}}$, one would generally expect the Jacobian $Du(\hat{\boldsymbol{x}}_{g})$ not to have purely imaginary eigenvalues.
2.1 Meta Tug-of-War Games
In general, a distributed algorithm that solves the QoS problem defined above may require significant coordination among players [6]. Remarkably, we identified a wide class of games where a simple distributed algorithm provably converges to an action profile with QoS guarantees. We term this class of games Meta-ToW, which naturally models common resource conflicts in a network of agents. We start by defining a single ToW game. As there is only a single game, we use $u_{n}(\boldsymbol{x})$ to denote the reward for player $n$.
@@ 2.1 Meta Tug-of-War Games
There could be multiple action profiles that meet the QoS guarantees. Since in ToW games the player’s action is a measure of “pull” (e.g., power), we prefer equilibrium points with smaller actions.
@@
The reward function also satisfies $u_{n}(\boldsymbol{g},\boldsymbol{x})=0$, if $x_{n}=0$ and $u_{n}(\boldsymbol{g},\boldsymbol{x})\geq 0$ for all $n,\boldsymbol{x}\in\mathcal{X}$ and $\boldsymbol{g}\in\mathcal{G}^{N}$.
3 Applications
In this section, we detail three useful special cases of Meta-ToW games and the QoS achievability problem.
3.1 Multi-Channel Power Control in Wireless Networks
In power control, the players are $N$ transmitter-receiver pairs, so transmitter $n$ wishes to transmit to receiver $n$. There are $K$ channels, and at each timestep, each player $n$ decides the channel $g_{n}$ to transmit over. For each player $n$, the action $x_{n}$ denotes the transmission power of player $n$ in channel $g_{n}$. The interference experienced by receiver $n$ is given by
$I_{n}(\boldsymbol{g},\boldsymbol{x})=\sum_{m\neq n:g_{m}=g_{n}}c_{m,n}x_{m},$
$\frac{\partial u_{n}(\boldsymbol{g},\boldsymbol{x})}{\partial x_{m}}=-\frac{c_{n,n}c_{m,n}x_{n}}{(N_{0}+I_{n}(\boldsymbol{g},\boldsymbol{x}))^{2}}<0$
for all $m\neq n$ such that $g_{m}=g_{n}$ and $\boldsymbol{x}\in\mathcal{X}^{o}$. There are multiple different assumptions on $\boldsymbol{\lambda}$ and the matrix $C=[c_{m,n}]$ in the literature [40, 15] which all result in the existence of a point that satisfies the QoS requirement. Taking the boundary $B_{n}$ to be large enough for each player $n$ would ensure that the QoS is achieved in the bounded region $0,B_{n}]$. Moreover, in the case of a single channel (or equivalently a single game in our framework) these assumptions also result in the uniqueness of the equilibrium point $\hat{\boldsymbol{x}}$ which satisfies $u_{n}(\boldsymbol{x})=\lambda_{n}$ for all players $n$. Trivially, this point is also the minimal equilibrium point.
In wireless networks, the locations of the devices, and thus the channel gains, are typically modeled as random [15]. Then, there is no need for Assumption 2 as the randomness in the channel gain matrix $C$ has a similar effect on the Jacobian’s eigenvalues.
3.2 Distributed Task Allocation with Cooperative Agents
Consider $N$ agents who collaborate to perform $K$ tasks [29, 31]. Each player decides which task to work on and how much effort they make on this task. The total utility obtained from a task depends on all the agents’ efforts working on that task, which can be thought of as the value of the progress rate of this task. This total utility is then divided between the agents working on that task, which represents their marginal contribution to the total value attributed to the progress in the task. This total utility is monotonically increasing and concave with respect to the total effort spent on that task. Progress rate typically increases sublinearly in effort or linearly at best if there is no overhead in collaboration. Then, the value of increased progress usually has diminishing returns. For example, computing a job in one week instead of two is a much more significant boost than computing a job in one minute instead of two. This reward structure incentivizes agents to choose tasks where their marginal contribution is more significant. This reward, which measures the local progress of an agent, can be naturally measured locally, as opposed to the global progress. Examples for applications are LLM agents, cloud computing, and multirobot manufacturing or construction.
3.3 Activation in Sensor Networks
Let the $N$ sensors be the players, where the probability of player $n$ being active is $p_{n}$. The action taken by player $n$ is $x_{n}=1-p_{n}$, i.e., the probability of being off. Clearly, $x_{n}$ is in $[0,1]$. We define the reward function for each player $n$ as
$u_{n}(\boldsymbol{x})=f(P_{n}(\boldsymbol{x}))-\alpha+\beta x_{n}.$
Here $P_{n}(\boldsymbol{x})$ is the probability that player $n$’s observation was successfully transmitted to the destination. This probability depends on the actions of all players as player $n$’s action affects the number of observations it makes, and others’ actions affect how many packets are relayed to the destination. $f(\cdot)$ is the monotonically increasing ‘value’ assigned to observations, which is typically concave since the marginal value of data decreases as the amount of data received increases. Finally, $\beta x_{n}$ is the ‘reward’ for consuming less energy thanks to being off with probability $x_{n}$. This can be thought of as $\beta-\beta p_{n}$, which is a shifted cost due to battery usage. Then $\alpha<0$ is just an offset term that ensures that $u_{n}(\boldsymbol{x})=0$ if $x_{n}=0$.
Unlike in previous applications, here the action $x_{n}$ cannot be thought of as ‘effort’ or the amount of resource used. The reward $u_{n}$ is not necessarily monotone in $x_{n}$ over $[0,1]$. Clearly $P_{n}(\boldsymbol{x})$ decreases and $\beta x_{n}$ increases with increasing $x_{n}$. Hence we can expect a peak at some point if we keep $\boldsymbol{x}_{-n}$ fixed and change $x_{n}$. The aim of the players in this game is to collect as much data as possible but they are limited by their energy consumption. The QoS requirement ensures that sufficient observations are obtained from all sensors. The “best” equilibrium in this example is one where they collect the most data, which is given by the minimal equilibrium.